QUESTION:

Tanry Industries ltd.is an SME that has 1.9 million shares outstanding. It just paid a dividend of GH¢ 0.05. It estimates that dividends will grow at 17.9% for two years and 11.6 % for another two years. The discount rate is 11.9% Thereafter, the dividend growth rate will be a constant of 7.5 %

(a) Compute the intrinsic share price.

(b) If the founding shareholders are willing to sell 47 percent of its shares to an outside investor, how much should they sell the shares?  

(c) If the market price of a share is GH¢ 2.17, what is the market value of the company?

(d) Why is there a difference between the market value and intrinsic value of the company?
Answer #1

А B с E H - 1 2 3 4 0 1 2 3 4 5 Year DPS Growth 5 6 0.0500 0.0590 0.0695 0.0776 0.0866 0.0931 17.90% 17.90% 11.60% 11.60% 7.5

A B с D E F G H 1 2 3 4 5 Oo 0.05 5 6 Year DPS Growth PV 1 2 3 4 =C5*(1+D6) =D5*(1+E6) =E5*(1+F6) =F5*(1+G6) =G5*(1+H6) 0.179